Greg is making a bike ramp in the shape of a right triangle. The base of the ramp is 4 inches more than twice its height, and the length of the incline is 4 inches less than three times its height. How high is the ramp? My question is in setting up the problem-do I have 2x+4 and 3x-4 for the formula c2=a2 + b2?
Answer by mangopeeler07(462)

x=height
2x+4=base
3x-4=incline

Therefore
x=a
2x+4=b
3x-4=c

These are the values to plug into a^2+b^2=c^2.